pmorgan:

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel? (The formula R⋅T=D , where R is the rate of speed, T is the time, and D is the distance can be used.) Enter your answer in the box.

Hoodmemes:

Ight so x the distance traveled by Car A and y the distance traveled by car B. The speed of car A can be represented as x/2.4 and for Car B is y/4.

Hoodmemes:

From that information we can write an equation relating the speed of both cars.

Hoodmemes:

dfx/2.4 = 22 + y/4,Since y = x, then:x/2.4 = 22 + x/4,Solving for x, we obtain:x= (blank) meters travel by both cars,So car (blank) has a speed of (blank) mph.